# Unraveling the Enigma of Double Descent: An In-depth Analysis through the Lens of Learned Feature Space

### v2

@@ 4.2 CIFAR-10
Figure 3: The phenomenon of double descent on ResNet18s trained on CIFAR-10 $(N=50000)$, under varying label noise of $p=[0\%,10\%,20\%]$ and the prediction accuracy $P$ when $p>0$. Optimized with SGD for 200 epochs and decreasing learning rate. The test error curve under $p=[0\%,10\%,20\%]$ performs the double descent phenomenon and the prediction accuracy $P$ of in-context noisy data learning showed a correlation to the generalization performance beyond context.
@@ 5 Conclusion
In this paper, we have studied the feature space of learned representations projected by various neural networks. By analyzing the behavior of the model when achieving a perfect fit on the training set, we observe that: as the model’s complexity and capacity increase, there is a clear and distinctive isolation of noisy data points within their original ‘class-mates’ in the learned feature space. This intriguing observation indicates that broader neural networks interpolate correct samples throughout the vicinity of noisy samples in the learned feature space, effectively ’isolating’ the impact of introduced label noise without explicit regularization techniques. We believe this implicit property plays a significant role in understanding how over-parameterization enhances generalization. We posit that the mechanism behind this phenomenon can be attributed to two key factors. First, the gradient descent optimization algorithm endeavors to strike a perfect balance among training data points to minimize the loss function, which can lead to the isolation of noisy data points. Second, the characteristics of high-dimensional space may facilitate a beneficial separation of noisy data points from clean data in the learned feature space.
@@
In the presence of noise, the learner cannot learn the full map by interpolation. Indeed, the bias-variance trade-off predicts a U-shaped performance curve with first the learner increasing out-of-sample performances from digesting examples and then losing performances by trying to fit the in-sample noisy data. However, a learner who makes good use of regularization can do better and learn how to account for uncertainty. Likely, the number of model parameters needed to learn both the underlying map and to handle the noise is larger than the number of parameters needed to learn only the map in the ideal case without noise (the interpolation point). If the model is a perfect learner with perfect regularization, we do not expect the manifestation of any double descent, but rather a gradual improvement in out-of-sample performances with the number of parameters until a saturation point is reached where performances cannot be improved without further examples. We expect that this saturation point should occur at a number of parameters larger than the interpolation point. Double descent is instead observed for imperfect learners that at first try to interpolate the noise and then start an intrinsic regularization via over-parameterization. Our particular model of the emergence of the double descent phenomenon is illustrated In Fig. 4.
While the present study provides valuable insights into the mechanisms leading to the occurrence of the double descent phenomenon, a comprehensive theoretical framework is yet to be developed. Further investigations are needed to unravel the intricate details and offer a deeper understanding of this intriguing phenomenon. Another aspect that warrants attention is the impact of the architectures of deep neural networks. Studies have demonstrated that neural networks exhibit diverse behaviors based on their depth and width, even when the total number of parameters is identical. It piques our curiosity to investigate whether this variation in network shape could influence the underlying mechanisms of the double descent phenomenon.
Reproducibility Statement   To facilitate reproducibility, we provide a detailed description of the neural architectures in Section 3 and full training details including exact settings of hyper-parameters in Appendix A. We plan to release the source codes of our experiments to the research community.
@@ A.1 Scaling of Model Parameterization
The scaling of model parameterization to the universal layer width unit $K$ is shown in Figure.5.
@@
‘ (a) SimpleFC (b) Five-layer CNNs (c) ResNet18s
Figure 5: Scaling of the number of parameters as model size with layer width unit k of the three neural architectures used in our experiments. We apply a logarithmic scale to the parameter counts $P$ of all neural architectures.
A.2 Experiment Setups
We describe the experiment setup along with hyper-parameters we used in the training of all neural architectures below:
  • Loss function: Cross-entropy loss is used to evaluate the model’s performance for all our experiments.
  • Data Normalization: MNIST images do not undergo any data normalization. CIFAR-10 images are subjected to normalization using the scale and shift values of $((0.5,0.5,0.5),(0.5,0.5,0.5))$.
  • Data Augmentation: No data-augmentation techniques are explicitly used in any of our experiments.
  • Regularization: No explicit regularization like weight decay or dropout was applied unless explicitly stated.
  • Initialization: we use the default initialization provided by PyTorch for all the layers.
  • Optimization: For all our experiments, we employ the Stochastic Gradient Descent (SGD) optimizer with zero momentum and a batch size of 128. The learning rate starts with an initial value of 0.05 and updates every 50 epochs. Below, we provide a precise description of our different learning rate schedules on SimpleFC and CNNs/ResNet18s:
    • Learning Rate Scheme for SimpleFC:
      $LR=\frac{0.05}{\sqrt{1+[epoch\div 50]}}$ (2)
    • Learning Rate Scheme for CNNs/ResNet18s:
      $lr=\frac{0.05}{\sqrt{1+[epoch*10]}}$ (3)
A.3 Additional Results
Here, we showcase additional experimental results that, due to page limitations, could not be included in the main context.
A.3.1 Model Performance in the Presence of High Label Noise
We provided here two additional experiment results when a high label noise ratio $p=40\%$ is introduced to the training dataset of MNIST and CIFAR-10. It is evident that noise has a substantial impact on the model’s classification accuracy, leading to a reduction. However, the correlation between $P$ and generalization performance remains consistency.
@@ A.3.1 Model Performance in the Presence of High Label Noise
‘ (a) SimpleFC on MNIST $(p=40\%)$ (b) CNN on CIFAR-10 $(p=40\%)$
@@
A.3.2 t-SNE visualization of Learned Representations
In addition to evaluating the value of $P$, we also utilize T-Distributed Stochastic Neighbor Embedding (t-SNE) to provide a visual representation of the learned features in the feature space. While this visualization may not accurately represent the intricate inherent structure of high-dimensional features, it aims to gain insights into the clustering and distribution of data points, thereby enhancing our understanding of the model’s internal representations.
The t-SNE visualization in Figure.7 offers evidence from a more intuitive standpoint: as layer width increases, the overlap between different classes diminishes, and noisy data points are more accurately assigned to their correct classes. The choice of layer width $k=[10,20,100]$ represents three different circumstances. When $k=10$, the bias-variance trade-off is reached at the bottom of the U-shaped segment of the test error curve. The training dataset is not fully interpolated and the learned representations are less organized. When $k=20$, the model is about to interpolate the whole training set, which includes the noisy labeled data. From the figure, it’s evident that at this stage, representations of some data points with noisy labels appear within classes they don’t belong to, which are likely to introduce interference and lead to misleading effects. In the case of an over-parameterized model, such as when $k=100$, this phenomenon significantly diminishes, and the majority of noisy labeled data points are encoded alongside their true classmates.
Figure 7: T-Distributed Stochastic Neighbor Embedding (T-SNE) Visualization of the learned representation vectors by trained SimpleFC of 1000 randomly selected data points from the training set of MNIST, at different layer widths of $k=[10,20,100]$. Circular points denote the clean data points, while asterisk points represent the data points with noisy labels (color-coded according to their original class). With an increase in layer width, the overlap between distinct classes diminishes, leading to a more precise allocation of noisy data to their respective correct classes, which indicates that over-parameterized models can segregate noise within the learned feature space.
When we examine the t-SNE visualization of learned representations in CNN trained on the CIFAR-10 dataset, despite the increased complexity of 2D distributions due to the higher dimensionality of CIFAR-10 images, we can still observe distinct class distributions and improved identification of noisy data points within these distributions as the model width expands. The t-SNE visualization Figure.9 of ResNets conveys the same statement.
Figure 8: T-Distributed Stochastic Neighbor Embedding (T-SNE) Visualization of the learned representation vectors by trained CNNs of 1000 randomly selected data points from the training set of CIFAR-10, at different layer widths of $k=[6,12,36]$. Circular points denote the clean data points, while asterisk points represent the data points with noisy labels (color-coded according to their original class). With an increase in layer width, the overlap between distinct classes diminishes, leading to a more precise allocation of noisy data to their respective correct classes, indicating that over-parameterized models can segregate noise within the learned feature space.
Figure 9: T-Distributed Stochastic Neighbor Embedding (T-SNE) Visualization of the learned representation vectors by trained ResNet18s of 1000 randomly selected data points from the training set of CIFAR-10, at different layer widths of $k=[6,12,36]$. Circular points denote the clean data points, while asterisk points represent the data points with noisy labels (color-coded according to their original class). With an increase in layer width, the overlap between distinct classes diminishes, leading to a more precise allocation of noisy data to their respective correct classes, which indicates that over-parameterized models can segregate noise within the learned feature space.

### v3

@@ 4 Experiments
Figure 3: The phenomenon of double descent on ResNet18s trained on CIFAR-10 $(N=50000)$, under varying label noise of $p=[0\%,10\%,20\%]$ and the prediction accuracy $P$ when $p>0$. Optimized with SGD for 200 epochs and decreasing learning rate. The test error curve under $p=[0\%,10\%,20\%]$ performs the double descent phenomenon. The k-NN prediction accuracy $P$ on the clean labels first increases and then decreases to 0 at the interpolation threshold. One additional purple line on predicting the noisy labels is introduced showcasing that the k-NN predicts all noisy data with their clean labels.
However, in Figure 3 of ResNet18s trained and examined on the sample task that w.r.t. increase in model width, $P$ increases before the test error peak, then decreases to zero at the interpolation threshold and remains consistent after since. By demonstrating a complementary prediction accuracy on the noisy labels rather than the true labels, we observe that for models larger the interpolation threshold, the k-NN predicts all noisy samples to their assigned noisy labels than the clean labels. In further conclusion, the learned representations of the noisy data largely correspond to the noisy information it is provided for over-parameterized models. We attribute this observation to the superior feature extraction ability of ResNets compared to CNNs by more convolutional layers and reduced training difficulty by residual connections.
In this experimental setup, the observation contradicts our hypothesis that over-parameterized models attain a better ability to ‘isolate’ noisy samples. However, it’s worth noting that while ResNets outperform CNNs of intermediate scale initially, the introduction of additional label noise leads to over-parameterized CNNs surpassing ResNets by 5% in test accuracy and even more in test losses. This comparison supports the conclusion that the k-NN prediction capability is linked to the generalization performance of deep learning models concerning their size. Throughout the studied neural architectures, that the injection of input features becomes more abstract in the hidden feature space, and more complex class feature patterns can be learned, even when noise is introduced. This observation leads to a further question: even though the benign over-parameterization phenomenon applies inside various network architectures, the issue of overfitting persists when comparing models of varying complexities.
@@ 5 Conclusion
In this paper, we have studied the feature space of learned representations projected by various neural networks. By analyzing the behaviour of the model when achieving a perfect fit on the training set, we observe that: as the model’s complexity and capacity increase, there is clear and distinctive isolation of noisy data points within their original ‘class-mates’ in the learned feature space. This intriguing observation indicates that broader neural networks interpolate correct samples throughout the vicinity of noisy samples in the learned feature space, effectively ‘isolating’ the impact of introduced label noise without explicit regularization techniques. We believe this implicit property plays a significant role in understanding how over-parameterization enhances generalization. We posit that the mechanism behind this phenomenon can be attributed to two key factors. First, the gradient descent optimization algorithm endeavours to strike a perfect balance among training data points to minimize the loss function, which can lead to the isolation of noisy data points. Second, the characteristics of high-dimensional space may facilitate a beneficial separation of noisy data points from clean data in the learned feature space.
@@
In the presence of noise, the learner cannot learn the full map by interpolation. Indeed, the bias-variance trade-off predicts a U-shaped performance curve with first the learner increasing out-of-sample performances and then losing performances by trying to fit the in-sample noisy data. However, a learner who makes good use of regularization can do better and learn how to account for uncertainty. Likely, the number of model parameters needed to learn both the underlying map and to handle the noise is larger than the number of parameters needed to learn only the map in the ideal case without noise (the interpolation threshold). If the model is a perfect learner with perfect regularization, we do not expect the manifestation of any double descent, but rather a gradual improvement in out-of-sample performances with the number of parameters until a saturation point is reached where performances cannot be improved without further examples. We expect that this saturation point should occur at a number of parameters larger than the interpolation point. Double descent is instead observed for imperfect learners who at first try to interpolate the noise and then start an intrinsic regularization via over-parameterization. Our particular model of the emergence of the double descent phenomenon is illustrated In Fig. 4.
While the present study provides valuable insights into the mechanisms leading to the occurrence of the double descent phenomenon, a comprehensive theoretical framework is yet to be developed. Further investigations are needed to unravel the intricate details and offer a deeper understanding of this intriguing phenomenon. Another aspect that warrants attention is the impact of the architectures of deep neural networks. Studies have demonstrated that neural networks exhibit diverse behaviours based on their depth and width, even when the total number of parameters is identical. It piques our curiosity to investigate whether this variation in network shape could influence the underlying mechanisms of the double descent phenomenon. Another potential avenue for future exploration could involve implicit sparsity, exhibited by feature representations in over-parameterized neural networks.
Reproducibility Statement To facilitate reproducibility, we provide a detailed description of the neural architectures and full training details including exact settings of hyper-parameters in Section 3. Our code is provided publicly available at https://github.com/Yufei-Gu-451/double_descent_inference.
@@ A.1 Scaling of Model Parameterization
The scaling of model parameterization to the universal layer width unit $K$ is shown in Figure.5. In our experiment results, we labelled the x-axis scale as the model width parameter $k$ instead of the scaling of parameterization. However, we believe this labelling does not impact our conclusion because of the monotonic relationship between these two factors.
@@
‘ (a) FCNN (b) Five-layer CNN (c) ResNet18
Figure 5: Scaling the number of parameters as model size with layer width unit k of the three neural architectures used in our experiments in the Methodology section. We apply a logarithmic scale to all neural architectures’ parameter counts $P$.
A.2 Model Performance in the Presence of High Label Noise
Here, we provided two additional experiment results when a high label noise ratio $p=40\%$ is introduced to the training dataset of MNIST and CIFAR-10 that, due to page limitations, could not be included in the main context. Noise has a substantial impact on the model’s classification accuracy, leading to a reduction in test accuracy. It is evident that the prediction accuracy curve $P$ exhibits significant instability when subjected to a high percentage of noisy samples, with correct samples becoming sparser in the learned feature space. However, the correlation between $P$ and generalization performance remains consistent.
@@ A.2 Model Performance in the Presence of High Label Noise
‘ (a) FCNN on MNIST $(p=40\%)$ (b) CNN on CIFAR-10 $(p=40\%)$